OpenStudy (anonymous):

describe the 3 situations when you can use the law of sines

OpenStudy (anonymous):



OpenStudy (anonymous):

when they give you 2 sides and 1 degrees and you have to find the other degrees as in example or the other way when they give you one side and 2 degrees and you have to find the other side than you have to use law of sins to find the side or degrees

OpenStudy (anonymous):

formula is like\[\frac{ \sin a }{ a }=\frac{ \sin b }{ b }=\frac{ \sin c }{ c }\]

OpenStudy (anonymous):

You can use it if you already have: Two Sides, One Angle (angle can be interior/exterior) (SAS) OR Two Angles, One Side (ASA)

OpenStudy (anonymous):

thanks guys.. but it says there are 3 situations

OpenStudy (anonymous):

isesnt that only

OpenStudy (anonymous):

two

OpenStudy (anonymous):

1)SAS interior angle 2)SAS exterior angle 3)ASA

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem :)